a current of i flows through a metallic wire of radius r and the free electrons in a drift with a velocity V D calculate the Drift velocity of free electron through the wires of the same material having double the radius when the same current flows through it
1
Expert's answer
2018-06-25T10:58:08-0400
r1=2*r2
I1=I2
n1=n2
v1=I1/e*n1*3.14*(r1)^2
v2=I2/e*n2*3.14*(r2)^2
v1/v2=4
That is, with an increase the radius of the wire in 2 times, the Drift velocity of free electron is reduced by 4 times.